# GLProtein: Global-and-Local Structure Aware Protein Representation Learning

### v1

@@ 3.4 Local Structure Information Modelling
Complementing global structural information, local protein information offers additional structure-related insights for protein language modelling. While the global structure information modelling component is designed to identify structural similarities across different proteins, the local structure information modelling component zooms in on the specific, intricate features of a protein’s internal structure, providing a more nuanced understanding. More specifically, in this part, we leverage the local structural details of proteins, including protein 3D distance encoding and substructure-based molecular encoding, to enhance the model’s ability to learn and interpret this local configuration effectively, as shown at the bottom of Figure 3. We will first detail the process of protein 3D distance encoding. Then, the integration of embedding of amino acid molecules will be illustrated.
3.4.1 Protein 3D Distance Encoding
In protein representation learning, the integration of protein 3D coordinates holds immense significance for comprehensively capturing proteins’ structural intricacies and functional behaviours. The 3D coordinates provide critical insights into how proteins fold and interact in three-dimensional space, influencing their stability, activity, and specificity liu2022generating ; peng2022pocket2mol ; su2023saprot . By incorporating 3D structural information into representation learning models, the model can capture proteins’ geometric and spatial characteristics, which are often overlooked by sequence-based approaches. We use AlphaFoldDB [^1] https://alphafold.ebi.ac.uk/ as the 3D protein database and integrate the protein 3D distance encoding ying2021transformers to represent protein 3D structural information to ensure rotational and translational invariance. However, proteins are complex macromolecules composed of numerous amino acids, each consisting of multiple atoms. Consequently, the atomic coordinate data for each protein is extensive. To address this challenge, we propose to take advantage of the alpha-carbon ($\alpha$-C) coordinates rather than the entire protein coordinates in protein representation learning. That is because the $\alpha$-C atoms are central to the protein backbone and provide a simplified yet informative representation of the protein’s 3D structure. By capturing the backbone conformation, $\alpha$-C coordinates effectively convey the protein’s overall shape and folding pattern, which are critical for understanding its function. Moreover, leveraging $\alpha$-C coordinates balances capturing essential structural information and maintaining computational efficiency.
Specifically, the coordinates of each $\alpha$-C are processed to represent the current position of the amino acid in 3D space. Then, we encode the Euclidean distance metric to reflect the spatial relation between any pair of amino acids in the 3D space. Mathematically, given each amino acid pair $(i,j)$, we first process their Euclidean distance with the Gaussian Basis Kernel function scholkopf1997comparing ,
$\phi^{k}_{(i,j)}=-\frac{1}{\sqrt{2\pi}|\sigma^{k}|}\exp(-\frac{1}{2}(\frac{\gamma_{(i,j)}||\mathbf{r_{i}}-\mathbf{r_{j}}||+\beta_{(i,j)}-\mu^{k}}{|\sigma^{k}|})^{2}),$ (4)
@@ 3.4.1 Protein 3D Distance Encoding
3.4.2 Substructure-based Molecular Encoding
Amino acid molecules, as more detailed information about protein localization, play a crucial role in protein representation learning, as they form the essential building blocks of proteins and provide the foundational data for understanding protein structure and function lieu2020amino ; lopez2024biochemistry . To learn the fine-grained amino acid structure information, we introduce substructure-based molecular encoding to leverage the inherent relationships between molecule motifs and substructural features in amino acid molecules. In practice, we utilize the mol2vec jaeger2018mol2vec to process and derive representations for all amino acid molecules to obtain fine-grained molecular structure information. Then we construct substructure-based molecular encoding as fine-grained protein sequence representation, which is the other part of local structure information modelling component. For protein $P$, we have
$\mathbf{E}_{a}(P)=\text{Concat}(\mathbf{e}_{x_{1}},\mathbf{e}_{x_{2}},\dots,\mathbf{e}_{x_{i}},\dots,\mathbf{e}_{x_{L}}),\text{where}~{}\mathbf{e}_{x_{i}}\in\mathbb{R}^{1\times d},$ (6)
$L$ is the length of the protein sequence, $\mathbf{e}_{x_{i}}$ is the $i$-th amino acid molecule embedding, and $d$ stands for the feature dimension of the amino acid molecule.
@@ 3.5 Model Training
3.5.1 Protein Decoder
As shown in Figure 3, to enhance the learning of protein representation, we introduce a protein decoder that enables each protein sequence representation to iteratively query and extract relevant information from its associated local structure, including protein 3D distance encoding and substructure-based molecular encoding. This process facilitates the integration of critical structural details, aiding in the accurate reconstruction of masked amino acids. Specifically, the decoder treats protein representation $\mathbf{E}_{p}$ as a query, while the substructure-based molecular encodings $\mathbf{E}_{a}$ are attended to as keys and values and protein 3D distance encoding $\Phi^{distance}$ is attended to as attention bias. Taking the $i$-th layer as an example, the inputs to the protein decoder include $\mathbf{E}_{p}^{i}$, $\Phi^{distance}$ and $\mathbf{E}_{a}$. The substructure-based molecular encoding $\mathbf{E}_{a}$ is firstly queries by $\mathbf{E}_{p}^{i}$ as the key and value:
$Q_{p}^{i}=\text{Norm}(\mathbf{E}_{p}^{i})W_{Q}^{i},~{}K_{a}^{i}=\text{Norm}(\mathbf{E}_{a})W_{K}^{i},~{}V_{a}^{i}=\text{Norm}(\mathbf{E}_{a})W_{V}^{i},$ (7)
where $W_{Q}^{i}$,$W_{K}^{i}$,$W_{V}^{i}$ are learnable matrices. Norm stands for the layer normalization ba2016layer .
Then, Attention vaswani2017attention is applied to $\{Q_{p}^{i},K_{a}^{i},V_{a}^{i}\}$, where the representation of protein sequence extract helpful, relevant information from the substructure-based molecular encoding. The obtained representation $o_{p}^{i}$ stores the helpful structure information for restoring missing amino aids. We then add up $o_{p}^{i}$ and $\mathbf{E}_{p}^{i}$ to integrate information, resulting in the representation $\hat{\mathbf{E}}_{p}^{i}$ as follows:
$o_{p}^{i}=\text{Attention}(Q_{p}^{i},K_{a}^{i},V_{a}^{i},\Phi^{distance}),~{}~{}\hat{\mathbf{E}}_{p}^{i}=\text{Norm}(\mathbf{E}_{p}^{i})+o_{p}^{i}.$ (8)
@@ 3.5.1 Protein Decoder
Figure 4: An example of amino acid contacts (top-L predictions for ProteinNet alquraishi2019proteinnet test example TBM-hard#T0912). Raw contact probabilities are shown below the diagonal, top L contacts are shown above the diagonal (blue: true positives, red: false positives, grey: ground-truth contacts).
By training with both protein 3D distance encoding and substructure-based molecular encoding, the model can not only go beyond simply memorizing sequence patterns but also actively decode the underlying structural and chemical principles that govern protein functionality.
3.5.2 Pre-training Objective
@@
Figure 5: Results on TAPE Benchmark encompass various evaluations. SS is a secondary structure task that is evaluated in CB315. We report medium- and long-range results using P@L/2 metrics in contact prediction task. In fluorescence and stability prediction tasks, we use Spearman’s $\rho$ metric for evaluation. We also provide a related table in Appendix 7.1.
@@ 4 Experiments
4.1 Dataset and Implementation Details
ProteinKG25  zhangontoprotein offers a comprehensive knowledge graph comprising approximately five million triplets. This graph includes nearly 600k proteins, 50k attribute terms and 31 relation terms. However, note that we only use the protein sequence data in this dataset. Our model does not use the attribute and relation terms extracted from Gene Ontology. We used modified pre-training data from KeAP zhou2023protein , which removed overlapping data between ProteinKG25 and downstream datasets. Additionally, we use AlphaFoldDB to obtain the protein 3D coordinate datasets.
We conducted some experiments and compared GLProtein with baselines regarding pre-training and inference time in contact prediction tasks, as shown in Appendix Table 4. GLProtein outperforms baselines in multiple downstream tasks with similar parameters. During pre-training, GLProtein is trained for 300k steps using a learning rate of 1e-5, weight decay of 0.01 over four GPUs (NVIDIA A6000, 48G Memory each). For the amino acid contact prediction and protein-protein interaction task, we randomly selected five random seeds to fine-tune our model and the baseline model separately and report the results. For full implementation details, refer to the provided code.
@@ 4.2 Downstream Tasks
4.2.1 Amino Acid Contact and Secondary Structure Prediction
@@
Overview of amino acid contact prediction. Amino acid contact prediction is a critical task in computational biology, aiming to identify pairs of amino acids within a protein that are in close spatial proximity. Given an input protein sequence, our model predicts whether pairs of amino acids from the same sequence are in contact. The model accomplishes this by generating a probability contact matrix for each input protein. We tested the model on the dataset collected and organized by ProteinNet alquraishi2019proteinnet and TAPE rao2019evaluating .
Overview of secondary structure prediction. Secondary structure is a fundamental aspect of computational biology, aiming to determine the local structures of protein segments. this task is a sequence-to-sequence task where each input protein is mapped to a type of local structure. We report accuracy on a per-amino acid basis on the CB513 dataset cuff1999evaluation .
@@
Baselines. We evaluate our model compared with ten baselines. Specifically, we employed variations of LSTM hochreiter1997long , ResNet he2016deep and Transformer vaswani2017attention proposed by the TAPE benchmark rao2019evaluating . ProtBert elnaggar2021prottrans is a BERT-like model pre-trained on UniRef100 suzek2007uniref ; suzek2015uniref . ESM-2 rives2021biological ; verkuil2022language ; hie2022high feature a transformer architecture pre-trained on the representative sequences from UniRef50 suzek2007uniref ; suzek2015uniref . OntoProtein zhangontoprotein and KeAP zhou2023protein are the most recent knowledge-based pre-training methodologies. SaProt su2023saprot is the most recent structure-based protein language model. LM-GVP wang2022lm and GearNet zhang2022protein are famous geometric methods for protein representation learning.
@@
Results. Table  1 shows the experimental results of amino acid contact prediction. Specifically, we notice that our model GLProtein consistently outperforms other models in short- ($126\leq seq<$ 12), medium- ($2412\leq seq<$ 24) and long-range ($seq>24$) contact predictions. Notably, our model demonstrates better performance compared to SaProt, which is also a structure-based language model. We also randomly selected a protein from the contact test dataset for visual analysis. As shown in Figure 4, the left is our GLProtein’s result of amino acid contacts. The right three are the contact maps of three baseline models, including KeAP, ESM-2 and ProtBERT. Figure 4 shows more visually that GLProtein’s prediction on the task of contact prediction is closer to labels, i.e., better performance on long-range contact prediction. For the secondary structure (SS-Q3 and SS-Q8), as shown in Figure 5, GLProtein outperforms other baselines in SS-Q8 task and shows competitive performance with ProtBERT, OntoProtein and KeAP in SS-Q3 task. Considering the approaches taken by Saprot, LM-GVP, and GearNet, which also incorporate protein structural information, the evident performance superiority of GLProtein over these methods indicates that it offers a more effective option for structure-based protein representation learning. We attribute the enhancements in performance achieved by GLProtein to its innovative integration of global and local structural information, which allows the pre-trained model to gain a deeper understanding of protein structure.
@@ 4.2.2 Protein-Protein Interaction
In our experiments, we focus on predicting 7 interaction types between protein pairs, namely reaction, binding, post-translational modifications, activation, inhibition, catalysis, and expression. The challenge of PPI prediction is approached as a multi-label classification problem. We conducted our experiments using three datasets: SHS27K chen2019multifaceted , SHS148K chen2019multifaceted and STRING lv2021learning . Both SHS27K and SHS148K are considered subsets of STRING, with proteins excluded if they have fewer than 50 amino acids or exhibit 40% or higher sequence identity. We followed OntoProtein’s setting to generate test sets and employed Breadth-First Search (BFS) and Depth-First Search (DFS) techniques across these datasets. The F1 score is utilized as the primary metric for evaluating performance.
Baselines. Following OntoProtein zhangontoprotein and KeAPzhou2023protein , we have expanded our baseline models to include four additional methods: DPPI hashemifar2018predicting , DNN-PPI li2018deep , PIPR chen2019multifaceted and GNN-PPI lv2021learning . These are incorporated alongside existing baselines such as ProtBert, ESM-2, OntoProtein, KeAP, SaProt, LM-GVP and GearNet, providing a comprehensive set of comparisons in our analysis.
@@
Results. As shown in Table 2, the results clearly indicate that our method consistently outperforms all other methods, including the structure-based protein language model SaProt, across all datasets and both BFS and DFS evaluation metrics. The observed decline in performance can be linked to the growing amount of fine-tuning data, transitioning from SHS27k to STRING, which diminished the influence of pre-training. We believe that the structural similarities between proteins identified during the pre-training step enable GLProtein to excel in the PPI task, resulting in its outstanding performance.
4.2.3 Homology Detection, Fluorescence and Stability Prediction

### v2

@@ 3.3 Local Structure Information Modelling
While the global structure information modelling component is designed to identify structural similarities across different proteins, the local structure information modelling component zooms in on the specific, intricate features of a protein’s internal structure, providing a more nuanced understanding. More specifically, in this part, we leverage the local structural details of proteins, including protein 3D distance encoding and substructure-based molecular encoding, to enhance the model’s ability to learn and interpret this local configuration effectively, as shown at the bottom of Figure 3.
Protein 3D Distance Encoding.
The 3D coordinates provide critical insights into how proteins fold and interact in three-dimensional space, influencing their stability, activity, and specificity Liu et al. (2022); Peng et al. (2022); Su et al. (2023). We use AlphaFoldDB [^1] https://alphafold.ebi.ac.uk/ as the 3D protein database and integrate the protein 3D distance encoding Ying et al. (2021) to represent protein 3D structural information to ensure rotational and translational invariance. We propose to take advantage of the alpha-carbon ($\alpha$-C) coordinates rather than the entire protein coordinates in protein representation learning. By capturing the backbone conformation, $\alpha$-C coordinates effectively convey the protein’s overall shape and folding pattern, which are critical for understanding its function. Moreover, leveraging $\alpha$-C coordinates balances capturing essential structural information and maintaining computational efficiency.
[…]
@@
Substructure-based Molecular Encoding.
As more detailed information about protein localisation, amino acid molecules play a crucial role in protein representation learning, as they form the essential building blocks of proteins and provide the foundational data for understanding protein structure and function Lieu et al. (2020); Lopez and Mohiuddin (2024). To learn the fine-grained amino acid structure information, we introduce substructure-based molecular encoding to leverage the inherent relationships between molecule motifs and substructural features in amino acid molecules. In practice, we utilize the mol2vec Jaeger et al. (2018) to process and derive representations for all amino acid molecules to obtain fine-grained molecular structure information. For protein $P$, we have
$\mathbf{E}_{a}(P)=\text{Concat}(\mathbf{e}_{x_{1}},\mathbf{e}_{x_{2}},\dots,\mathbf{e}_{x_{i}},\dots,\mathbf{e}_{x_{L}}),$
where $\mathbf{e}_{x_{i}}\in\mathbb{R}^{1\times d}$, $L$ is the length of the protein sequence, $\mathbf{e}_{x_{i}}$ is the $i$-th amino acid molecule embedding, and $d$ stands for the feature dimension of the amino acid molecule.
@@ 3.4 Model Training
Protein Decoder.
As shown in Figure 3, the decoder treats protein representation $\mathbf{E}_{p}$ as a query, while the substructure-based molecular encodings $\mathbf{E}_{a}$ are attended to as keys and values and protein 3D distance encoding $\Phi^{distance}$ is attended to as attention bias. Taking the $i$-th layer as an example, the inputs to the protein decoder include $\mathbf{E}_{p}^{i}$, $\Phi^{distance}$ and $\mathbf{E}_{a}$. The substructure-based molecular encoding $\mathbf{E}_{a}$ is firstly queries by $\mathbf{E}_{p}^{i}$ as the key and value:
$\begin{split}Q_{p}^{i}=\text{Norm}(\mathbf{E}_{p}^{i})W_{Q}^{i},\\ K_{a}^{i}=\text{Norm}(\mathbf{E}_{a})W_{K}^{i},\\ V_{a}^{i}=\text{Norm}(\mathbf{E}_{a})W_{V}^{i},\end{split}$
where $W_{Q}^{i}$$W_{K}^{i}$$W_{V}^{i}$ are learnable matrices. Norm stands for the layer normalization Ba et al. (2016).
Then, Attention Vaswani et al. (2017) is applied to $\{Q_{p}^{i},K_{a}^{i},V_{a}^{i}\}$, where the representation of protein sequence extracts helpful, relevant information from the substructure-based molecular encoding. The obtained representation $o_{p}^{i}$ stores the helpful structure information for restoring missing amino aids. We then add up $o_{p}^{i}$ and $\mathbf{E}_{p}^{i}$ to integrate information, resulting in the representation $\hat{\mathbf{E}}_{p}^{i}$ as follows:
$\begin{split}&o_{p}^{i}=\text{Attention}(Q_{p}^{i},K_{a}^{i},V_{a}^{i},\Phi^{distance}),\\ &\hat{\mathbf{E}}_{p}^{i}=\text{Norm}(\mathbf{E}_{p}^{i})+o_{p}^{i}.\end{split}$
@@
Figure 4: An example of amino acid contacts (top-L predictions for ProteinNet AlQuraishi (2019) test example TBM-hard#T0912). Raw contact probabilities are shown below the diagonal, top L contacts are shown above the diagonal (blue: true positives, red: false positives, grey: ground-truth contacts).
Pre-training Objective.
@@
Figure 5: Results on TAPE Benchmark encompass various evaluations. $S$ is a secondary structure task that is evaluated in CB315. We report medium- and long-range results using P@L/2 metrics in contact prediction task. In fluorescence and stability prediction tasks, we use Spearman’s $\rho$ metric for evaluation. We also provide a related table in Appendix A.1.
@@ 4 Experiments
Pretraining Datasets.
Swiss-Prot Boeckmann et al. (2003) offers a comprehensive and manually curated protein sequence database that includes nearly 600k protein sequences. We use it as pertaining dataset. Additionally, we use AlphaFoldDB to obtain the protein 3D coordinate datasets.
Implementation Details.
We conducted some experiments and compared GLProtein with baselines regarding pre-training and inference time in contact prediction tasks, as shown in Appendix Table 5. GLProtein outperforms baselines in multiple downstream tasks with similar parameters. During pre-training, GLProtein is trained for 300k steps using a learning rate of 1e-5, weight decay of 0.01 over four GPUs (NVIDIA A6000, 48G Memory each). We uniformly fine-tuned all downstream tasks without structural information to ensure fair and unbiased comparisons. For the amino acid contact prediction and protein-protein interaction task, we randomly selected five random seeds to fine-tune our model and the baseline model separately and report the results.
@@ 4.1 Downstream Tasks
4.1.1 Amino Acid Contact Prediction
@@
Overview. Amino acid contact prediction is a critical task in computational biology, aiming to identify pairs of amino acids within a protein that are in close spatial proximity. Given an input protein sequence, our model predicts whether pairs of amino acids from the same sequence are in contact. The model accomplishes this by generating a probability contact matrix for each input protein. We tested the model on the dataset collected and organized by ProteinNet AlQuraishi (2019) and TAPE Rao et al. (2019).
@@
Baselines. We evaluate our model compared with ten baselines. Specifically, we employed variations of LSTM Hochreiter and Schmidhuber (1997), ResNet He et al. (2016) and Transformer Vaswani et al. (2017) proposed by the TAPE benchmark Rao et al. (2019). ProtBert Elnaggar et al. (2021) is a BERT-like model pre-trained on UniRef100 Suzek et al. (2007, 2015). ESM-2 Rives et al. (2021); Verkuil et al. (2022); Hie et al. (2022) feature a transformer architecture pre-trained on the representative sequences from UniRef50 Suzek et al. (2007, 2015). OntoProtein Zhang et al. and KeAP Zhou et al. (2023a) are the most recent knowledge-based pre-training methodologies. SaProt Su et al. (2023) is the most recent structure-based protein language model. LM-GVP Wang et al. (2022) and GearNet Zhang et al. (2022) are famous geometric methods for protein representation learning. We uniformly fine-tuned all downstream tasks without structural information to ensure fair and unbiased comparisons. Since all structural tokens are masked, residual information still exists; we substitute Foldseek structure tokens with "#" when fine-tuning SaProt.
@@
Results. Table  1 shows the experimental results of amino acid contact prediction. Specifically, we notice that our model GLProtein consistently outperforms other models in short- ($6\leq seq<12$), medium- ($12\leq seq<24$) and long-range ($seq>24$) contact predictions. Notably, our model demonstrates better performance compared to SaProt, which is also a structure-based language model. We also randomly selected a protein from the contact test dataset for visual analysis. As shown in Figure 4, the left is our GLProtein’s result of amino acid contacts. The right three are the contact maps of three baseline models, including KeAP, ESM-2 and ProtBERT. Figure 4 shows more visually that GLProtein’s prediction on the task of contact prediction is closer to labels, i.e., better performance on long-range contact prediction. We attribute the enhancements in performance achieved by GLProtein to its innovative integration of global and local structural information, which allows the pre-trained model to gain a deeper understanding of protein structure. More results can be found in the Appendix A.1.
@@ 4.1.2 Protein-Protein Interaction
In our experiments, we focus on predicting 7 interaction types between protein pairs, namely reaction, binding, post-translational modifications, activation, inhibition, catalysis, and expression. The challenge of PPI prediction is approached as a multi-label classification problem. We conducted our experiments using three datasets: SHS27K Chen et al. (2019), SHS148K Chen et al. (2019) and STRING Lv et al. (2021). Both SHS27K and SHS148K are considered subsets of STRING, with proteins excluded if they have fewer than 50 amino acids or exhibit 40% or higher sequence identity. We followed OntoProtein’s setting to generate test sets and employed Breadth-First Search (BFS) and Depth-First Search (DFS) techniques across these datasets. The F1 score is utilized as the primary metric for evaluating performance.
Baselines. Following OntoProtein Zhang et al. and KeAPZhou et al. (2023a), we have expanded our baseline models to include four additional methods: DPPI Hashemifar et al. (2018), DNN-PPI Li et al. (2018), PIPR Chen et al. (2019) and GNN-PPI Lv et al. (2021). These are incorporated alongside existing baselines such as ProtBert Elnaggar et al. (2021), ESM-2 Beal (2015), OntoProtein Zhang et al. , KeAP Zhou et al. (2023a), SaProt Su et al. (2023), LM-GVP Wang et al. (2022), GearNet Zhang et al. (2022), DeepInter Lin et al. (2023), MAPE-PPI Wu et al. (2024), ProLLM Jin et al. (2024) and ESM-C Hayes et al. (2025), providing a comprehensive set of comparisons in our analysis.
@@
Results. As shown in Table 2, the results clearly indicate that our method consistently outperforms all other methods, including the structure-based protein language model such as SaProt and multimodal protein language model such as ESM-C, across all datasets and both BFS and DFS evaluation metrics. The observed decline in performance can be linked to the growing amount of fine-tuning data, transitioning from SHS27k to STRING, which diminished the influence of pre-training. We believe that the structural similarities between proteins identified during the pre-training step enable GLProtein to excel in the PPI task, resulting in its outstanding performance.
5 Conclusion and Future Work
In this work, we propose GLProtein, a general protein language model with global and local protein structure information. GLProtien outperforms the previous protein representation learning model on most downstream applications, demonstrating the performance superiority of GLProtein. In the future, we aim to further enhance GLProtein’s capabilities by exploring novel avenues for incorporating multi-modal data sources, refining the model’s interpretability, and extending its applicability to a wider array of biological contexts.
6 Limitations
7 Acknowledgements
The research described in this paper has been partially supported by the National Natural Science Foundation of China (project no. 62102335 and 62372314), General Research Funds from the Hong Kong Research Grants Council (project no. PolyU 15207322, 15200023, 15206024, and 15224524), internal research funds from Hong Kong Polytechnic University (project no. P0042693, P0048625, and P0051361). This work was supported by computational resources provided by The Centre for Large AI Models (CLAIM) of The Hong Kong Polytechnic University.
Appendix A Appendix
Table of Content:
  • Appendix A.1 Results on TAPE Benchmark
  • Appendix A.2 Secondary structure prediction
  • Appendix A.3 Homology Detection, Fluorescence and Stability Prediction
  • Appendix A.4 Protein Function Prediction
  • Appendix A.5 Protein-Protein Binding Affinity Estimation
  • Appendix A.6 Ablation Study
  • Appendix A.7 Parameter Sensitivity Study
  • Appendix A.8 Visualization of Protein Representations
  • Appendix A.9 Time Complexity Analysis
  • Appendix A.10 Hyper-parameters for Fine-tuning
A.1 Results on TAPE Benchmark
In addition to the Figure version, we also provide results on TAPE benchmark in a tabular version. As shown in Table 6, our model GLProtein performs competitive performance on many tasks, especially on the contact prediction and stability prediction tasks.
A.2 Secondary structure prediction
Overview. Secondary structure is a fundamental aspect of computational biology, aiming to determine the local structures of protein segments. this task is a sequence-to-sequence task where each input protein is mapped to a type of local structure. We report accuracy on a per-amino acid basis on the CB513 dataset Cuff and Barton (1999). Baselines. We evaluate our model compared with ten baselines. Specifically, we employed variations of LSTM Hochreiter and Schmidhuber (1997), ResNet He et al. (2016) and Transformer Vaswani et al. (2017) proposed by the TAPE benchmark Rao et al. (2019). ProtBert Elnaggar et al. (2021) is a BERT-like model pre-trained on UniRef100 Suzek et al. (2007, 2015). ESM-2 Rives et al. (2021); Verkuil et al. (2022); Hie et al. (2022) feature a transformer architecture pre-trained on the representative sequences from UniRef50 Suzek et al. (2007, 2015). OntoProtein Zhang et al. and KeAP Zhou et al. (2023a) are the most recent knowledge-based pre-training methodologies. SaProt Su et al. (2023) is the most recent structure-based protein language model. LM-GVP Wang et al. (2022) and GearNet Zhang et al. (2022) are famous geometric methods for protein representation learning.
Results. For the secondary structure (SS-Q3 and SS-Q8), as shown in Figure 5, GLProtein outperforms other baselines in SS-Q8 task and shows competitive performance with ProtBERT, OntoProtein and KeAP in SS-Q3 task. Considering the approaches taken by Saprot, LM-GVP, and GearNet, which also incorporate protein structural information, the evident performance superiority of GLProtein over these methods indicates that it offers a more effective option for structure-based protein representation learning. We attribute the enhancements in performance achieved by GLProtein to its innovative integration of global and local structural information, which allows the pre-trained model to gain a deeper understanding of protein structure.
A.3 Homology Detection, Fluorescence and Stability Prediction